# Optimal Algorithm for Multiple-Processor Multitask Scheduling

### v1

@@ 4.1 Average response time bound
$\mathbbm{E}{[F_{\rho}^{\mathsf{NP-SRPT}}]}\leq\mathbbm{E}{[F_{\rho}^{\mathrm{SRPT}_{1,N}}]}+O\Big{(}\log\frac{1}{1-\rho}\Big{)}\cdot\mathbbm{E}[\eta].$ (4)
Proof: Similar as the techniques in [20, 29], we relate the response time of the tagged job with an appropriate busy period.
Consider a tagged job with workload $x$, arriving time $r_{x}$ and completion time $C_{x}$. The computing resources of $N$ servers must be spent on the following types of job during $[r_{x},C_{x}]$:
@@
    The system may be processing jobs with rank larger than $x$, or some machines are idle, while the tagged job is in service, because the number of jobs alive is smaller than $N$. We use $\mathsf{W}_{\mathrm{waste}}(r_{x})$ to represent the amount of such resources, then
@@
    The servers may be dealing with system relevant $x$-work, which is of a higher priority than the tagged job. According to Lemma 8 and Lemma 9,
    $\mathsf{Rel}^{\mathsf{NP-SRPT}}_{\leq x}(r_{x})$ (6)
    $\leq$ $\mathsf{Rel}^{\mathrm{SRPT}_{1,N}}_{\leq x}(r_{x})+(N-1)\cdot(x+\eta)$
    $=$ $\mathsf{W}^{\mathrm{SRPT}_{1,N}}_{\leq x}(r_{x})+\mathsf{V}^{\mathrm{SRPT}_{1,N}}_{\leq x}(r_{x})+(N-1)\cdot(x+\eta)$
    $\leq$ $\mathsf{W}^{\mathrm{SRPT}_{1,N}}_{\leq x}(r_{x})+N\cdot(x+\eta).$ (7)
@@
[…]
@@
    Tagged job itself. The amount of resources is equal to $x$, the size of the tagged job.
Hence $\mathsf{T}^{\mathsf{NP-SRPT}}_{x}$, the response time of the tagged job, is no more than the length of a busy period of a single server system with speed $N$, which starts at time $r_{x}$ and has a initial workload of
$\mathsf{W}_{\mathrm{waste}}(r_{x})+\mathsf{Rel}^{\mathsf{NP-SRPT}}_{\leq x}(r_{x})+x.$
Combining with the aforementioned analysis, formally we have
$\mathsf{T}^{\mathsf{NP-SRPT}}_{x}$
$\leq_{st}$ $\mathsf{B}^{(\rho({x}))}\Big{(}\mathsf{W}_{\mathrm{waste}}(r_{x})+\mathsf{Rel}^{\mathsf{NP-SRPT}}_{\leq x}(r_{x})+x\Big{)}$
$\overset{(a)}{\leq}$ $\mathsf{B}^{(\rho({x}))}\Big{(}2Nx+N\eta+\mathsf{W}^{\mathrm{SRPT}_{1,N}}_{\leq x}(r_{x})\Big{)}$
=$\overset{(b)}{=}$ $\underbrace{\mathsf{B}^{(\rho({x}))}\Big{(}2Nx+N\eta\Big{)}}_{\Sigma_{1}}+\underbrace{\mathsf{B}^{(\rho({x}))}\Big{(}\mathsf{W}^{\mathrm{SRPT}_{1,N}}_{\leq x}(r_{x})\Big{)}}_{\Sigma_{2}}.$
In $(a)$ we use the upper bounds established in (5) and (7); $(b)$ follows from the additivity of busy period in M/G/1;
Note that the average response time under SRPT in a single server system is lower bounded as
@@
where the first equality holds due to the Poission Arrivals See Time Average (PASTA) property [30]. Note that
$\mathbbm{E}[\Sigma_{1}]$ $=O\Big{(}\mathbbm{E}\Big{(}\mathsf{B}^{(\rho(x))}(\eta+x)\Big{)}\Big{)}=O\Big{(}\mathbbm{E}\Big{[}\frac{\eta+x}{1-\rho(x)}\Big{]}\Big{)}$
In addition,
$\leq$ $\frac{1}{1-\rho(\xi)}+\frac{1}{\xi}\cdot\int_{\xi}^{\infty}\frac{xf(x)}{1-\rho(x)}\mathrm{d}x,$ (9)
where $\xi$ satisfies that $\rho(\xi)=\rho/2$. Note that
hence we have $\xi\geq\mathbbm{E}[p_{i}]/2$. Then the right hand side of (4.1) can be further bounded as
$=$ $2+\frac{2}{\mathbbm{E}[p_{i}]}\cdot\log\frac{1}{1-\rho}.$
Therefore for any input instance, the average response time under $\mathsf{NP}$-$\mathsf{SRPT}$, is no more than,
$\mathbbm{E}{[F_{\rho}^{\mathsf{NP-SRPT}}]}$ $=\mathbbm{E}_{x,r_{x}}{[\mathsf{T}^{\mathsf{NP-SRPT}}_{x}]}=\mathbbm{E}_{x,r_{x}}[\Sigma_{1}]+\mathbbm{E}_{x,r_{x}}[\Sigma_{2}]$
@@
The proof is complete. $\Box$
4.2 Optimality of $\mathsf{NP}$-$\mathsf{SRPT}$ with finite task size
@@
$\mathbbm{E}{[F^{\pi^{*}}_{\rho}]}\geq\mathbbm{E}{[F^{\mathrm{SRPT}_{1,N}}_{\rho}]}.$ (11)
@@
$\lim_{n\rightarrow\infty}\frac{\mathbbm{E}{[F_{\rho}^{\mathsf{NP-SRPT}}]}-\mathbbm{E}[F^{\mathrm{SRPT}_{1,N}}_{\rho}]}{\mathbbm{E}[F^{\mathrm{SRPT}_{1,N}}_{\rho}]}=0,$ (12)
which holds according to Lemma 17 and inequality (4.1)-(11).
5 Optimality of $\mathsf{NP}$-$\mathsf{SRPT}$ Beyond Finite Task Size
Up to this point, we have focused on job size distributions with finite task, which is rather restrictive. It is natural to consider relaxations of this assumption. In this section, we turn to other classes of job size distributions with unbounded support on task workload. These results provide complement to our developments about the theory of the asymptotic optimality of $\mathsf{NP}$-$\mathsf{SRPT}$
5.1 Warm up–random number of tasks
@@
Proof: Consider an increasing convex function $g(x)$, by Jensen’s inequality, we have
@@
where $D(s)=\{s>0|\mathbbm{E}[e^{sp_{i}}]<\infty\}$. The proof is complete. $\Box$
5.2 Optimality in M/M/1
@@
$\mathsf{NP}$-$\mathsf{SRPT}$ is asymptotic optimal in M/M/1.
@@
For M/M/1 model and any work-conserving algorithm, let $n_{\mathrm{busy}}$ be the number of arrivals in a busy period, then
@@
Proof of Theorem 19: Firstly, it has been shown in [2] that, the average response time under SRPT in M/M/1 satisfies
@@
$\mathbbm{E}{[F_{\rho}^{\mathsf{NP-SRPT}}]}-\mathbbm{E}{[F_{\rho}^{\mathrm{SRPT}_{1,N}}]}\leq O\Big{(}\log^{2}\frac{1}{1-\rho}\Big{)},$
which is a lower order term compared to $\mathbbm{E}[F^{\mathrm{SRPT-1}}_{\rho}]$. $\Box$
5.3 Beyond Exponential Job Size Distribution
In the study of $\mathsf{NP}$-$\mathsf{SRPT}$, it is crucial to analyze the expected length of busy periods, which could help bound $\mathbbm{E}[\eta]$. Notably, previous works have established insights into the analysis of busy periods in M/M/c queuing systems[32, 33]. In this section, we extend the analysis in [32] to M/G/N and prove an upper bound on the expected length of the busy period under $\mathsf{NP}$-$\mathsf{SRPT}$, with the following assumption.
@@ 5.3 Beyond Exponential Job Size Distribution
Proof: To analyze the busy periods under the $\mathsf{NP}$-$\mathsf{SRPT}$ policy, we classify two types of busy periods in an $N$-server system:
  • Full busy period. A time interval during which all $k\leq N$ servers are occupied. It starts when a new arrival finds $k-1$ customers already in the system and ends at the first departure epoch when exactly $k-1$ customers remain in the system. The length of such a busy period is denoted by $T_{k}$.
@@
$=$ (departure $\mathbbm{P}(\text{departure first})\cdot\mathbbm{E}[\text{Time tofirst departure}|\text{departure first}]$
+(arrival $+\mathbbm{P}(\text{arrival first})\cdot\mathbbm{E}[T_{k+1}|\text{start with}\{\mathsf{J}^{\prime}_{i}\}_{i=1}^{k}\cup\{\mathsf{J}_{k+1}\}]$
@@
Since we consider an $M/G$ system, where job arrivals follow a Poisson process, the probability that an arrival occurs before a departure is:
(arrival $\mathbbm{P}(\text{arrival first})$
$=$ $\int_{0}^{\infty}\mathbbm{P}(\text{minimum remaining workload of jobs under processing}\geq s)\cdot\lambda e^{-\lambda s}\mathrm{d}s$
@@
Using (5.3), we derive
$L_{k}(p_{\max})\leq\frac{1}{\mu}+\Big{(}1-\frac{1}{2}e^{-\frac{2r_{\max}^{2}}{N}}\Big{)}(L_{k+1}(p_{\max})+L_{k}(p_{\max}))$
$\Rightarrow$ $L_{k}(p_{\max})\leq\frac{2e^{\frac{2r_{\max}^{2}}{N}}}{\mu}+\Big{(}2e^{\frac{2r_{\max}^{2}}{N}}-1\Big{)}L_{k+1}(p_{\max}).$
@@
Let $\mathbb{E}[\mathsf{B}^{(N)}_{\mathsf{NP-SRPT}}]$ denote the expected duration of a partial busy period, defined as the time from the first job arrival in an empty $N$-server system until the system becomes empty again under the $\mathsf{NP}$-$\mathsf{SRPT}$ policy. Then
$\mathbbm{E}[\mathsf{B}^{(N)}_{\mathsf{NP-SRPT}}]\leq$ $\mathbbm{E}_{p_{\max}}\Big{[}\sum_{k=1}^{N}{L_{k}(p_{\max})}\Big{]}$
$=$ $O\Big{(}\frac{\log\mathbbm{E}[\mathsf{B}^{(N)}_{\mathsf{NP-SRPT}}]}{1-\rho}\Big{)}.$
@@
$\mathbbm{E}[\mathsf{B}^{(N)}_{\mathsf{NP-SRPT}}]$ $=O\Big{(}\frac{1}{(1-\rho)^{1+\epsilon}}\Big{)},\forall\epsilon>0.$
This completes the proof, following the arguments in Proposition 18. $\Box$
@@
Let $f$ be a positive function defined in $[0,\infty)$, the upper Matuszewska index is defined as the infimum of $\alpha$ for which there exists a constant $C=C(\alpha)$ such that for each $\bar{\lambda}>1$,
$\lim_{x\rightarrow\infty}\frac{f(\lambda x)}{f(x)}\leq C\lambda^{\alpha},$
@@
In an M/G/1 queue, if the upper Matuszewska index of the job size distribution is less than $-2$, then
@@
The average response time under $\mathsf{NP}$-$\mathsf{SRPT}$ is asymptotic optimal in M/GI/N under Assumption 21, if upper Matuszewska index of job size distribution is less than $-2$ and
@@
6 Experimental Results

### v2

@@ IV-A Average response time bound
$\mathbbm{E}{[F_{\rho}^{\mathsf{NP\text{-}SRPT}}]}\leq\mathbbm{E}{[F_{\rho}^{\mathsf{SRPT}_{1,N}}]}+O\Big(\log\frac{1}{1-\rho}\Big)\cdot\mathbbm{E}[\eta].$ (4)
Proof.
In this proof, we focus on the careful application of the technique in [12] to the complex $\mathsf{NP\text{-}SRPT}$ setting, correctly accounting for the additional delays caused by non-preemptivity and the multi-processor environment.
Proof Sketch: Similar as [12], the core idea is that the tagged job’s response time is determined by the time required for the $N$ servers to process a total volume of work, which includes the job’s own workload plus all other work that takes priority. By establishing upper bounds on each component and leveraging $\mathsf{M}/\mathsf{G}/1$ busy period properties, we derive a bound on the response time. Finally, we apply the PASTA property to relate this to the average response time of an ideal $\mathsf{SRPT}$ system.
Workload Decomposition for a Tagged Job and Bounding the Workload Components
Consider a tagged job with workload $x$, arriving time $r_{x}$ and completion time $C^{\mathsf{NP\text{-}SRPT}}_{x}$. Its response time is $\mathsf{T}^{\mathsf{NP\text{-}SRPT}}_{x}=C^{\mathsf{NP\text{-}SRPT}}_{x}-r_{x}$. This duration is the time for the $N$ servers to clear a specific volume of work. we categorize this work into four components processed during during $r_{x},C^{\mathsf{NP\text{-}SRPT}}_{x}]$:
@@
    Wasted service capacity. The system may be processing jobs with rank larger than $x$, or some machines are idle, while the tagged job is in service, because the number of jobs alive is smaller than $N$. We use $\mathsf{W}_{\mathrm{waste}}(r_{x})$ to represent the amount of such resources, then
@@
    Relevant work at arrivals. The servers may be dealing with system relevant $x$-work, which is of a higher priority than the tagged job. According to Lemma 1 and Lemma 2,
    $\mathsf{Rel}^{\mathsf{NP\text{-}SRPT}}_{\leq x}(r_{x})$ (6)
    $\leq$ $\mathsf{Rel}^{\mathsf{SRPT}_{1,N}}_{\leq x}(r_{x})+(N-1)\cdot(x+\eta)$
    $=$ $\mathsf{W}^{\mathsf{SRPT}_{1,N}}_{\leq x}(r_{x})+\mathsf{V}^{\mathsf{SRPT}_{1,N}}_{\leq x}(r_{x})+(N-1)\cdot(x+\eta)$
    $\leq$ $\mathsf{W}^{\mathsf{SRPT}_{1,N}}_{\leq x}(r_{x})+N\cdot(x+\eta).$ (7)
@@
    New relevant work. A newly arriving job is admitted to the system during $r_{x},C_{x}]$, only if its size is no more than $x$. Hence we only consider relevant load $\rho(x)=\lambda\cdot\int_{0}^{x}{tf(t)dt}$.
@@
    Tagged job’s workload. The amount of resources is equal to $x$, the size of the tagged job.
Constructing the Busy Period Bound
The response time $\mathsf{T}^{\mathsf{NP\text{-}SRPT}}_{x}$ of the tagged job $x$, is stochastically bounded by the length of a busy period in a single server system with speed $N$, driven by the relevant arrival load $\rho(x)$, starting with the sum of the tagged job’s work, wasted capacity, and pre-existing higher-priority work, i.e.,
$\mathsf{W}_{\mathrm{waste}}(r_{x})+\mathsf{Rel}^{\mathsf{NP\text{-}SRPT}}_{\leq x}(r_{x})+x.$
Formally, we have
$\mathsf{T}^{\mathsf{NP\text{-}SRPT}}_{x}$
$\leq_{st}$ $\mathsf{B}^{(\rho({x}))}\Big(\mathsf{W}_{\mathrm{waste}}(r_{x})+\mathsf{Rel}^{\mathsf{NP\text{-}SRPT}}_{\leq x}(r_{x})+x\Big)$
$\overset{(a)}{\leq}$ $\mathsf{B}^{(\rho({x}))}\Big(2Nx+N\eta+\mathsf{W}^{\mathsf{SRPT}_{1,N}}_{\leq x}(r_{x})\Big)$
$\overset{(b)}{=}$ $\underbrace{\mathsf{B}^{(\rho({x}))}\Big(2Nx+N\eta\Big)}_{\Sigma_{1}}+\underbrace{\mathsf{B}^{(\rho({x}))}\Big(\mathsf{W}^{\mathsf{SRPT}_{1,N}}_{\leq x}(r_{x})\Big)}_{\Sigma_{2}}.$
In $(a)$ we use the upper bounds established in (5) and (7); $(b)$ follows from the additivity of busy period in $\mathsf{M}$$\mathsf{G}$/11.
Asymptotic Analysis of the Bound
We analyze the expectations of $\Sigma_{1}$ and $\Sigma_{2}$. The term $\Sigma_{1}$ represents delay due to non-preemptivity and multi-server setup:
$\mathbbm{E}_{x,r_{x}}[\Sigma_{1}]$ $=O\Big(\mathbbm{E}\Big(\mathsf{B}^{(\rho(x))}(\eta+x)\Big)\Big)=O\Big(\mathbbm{E}\Big[\frac{\eta+x}{1-\rho(x)}\Big]\Big)$
Splitting the integral at $\xi$ such that $\rho(\xi)=\rho/2$:
$\leq$ $\frac{1}{1-\rho(\xi)}+\frac{1}{\xi}\cdot\int_{\xi}^{\infty}\frac{xf(x)}{1-\rho(x)}\mathrm{d}x.$ (9)
Note that
we have $\xi\geq\mathbbm{E}[p_{i}]/2$. Substituting into (IV-A):
$=$ $2+\frac{2}{\mathbbm{E}[p_{i}]}\cdot\log\frac{1}{1-\rho}.$
Thus,
$\mathbbm{E}_{x,r_{x}}{[\Sigma_{1}]}=O\Big(\log\frac{1}{1-\rho}\Big)\cdot\mathbbm{E}[\eta]$
The term $\Sigma_{2}$ is the busy period generated by higher-priority work in $\mathsf{SRPT}$ as seen by an arriving job. By the PASTA property [48]:
@@
Putting the Pieces Together
Combining the expectations of $\Sigma_{1}$ and $\Sigma_{2}$,
$\mathbbm{E}{[F_{\rho}^{\mathsf{NP\text{-}SRPT}}]}$ $=\mathbbm{E}_{x,r_{x}}{[\mathsf{T}^{\mathsf{NP\text{-}SRPT}}_{x}]}=\mathbbm{E}_{x,r_{x}}[\Sigma_{1}]+\mathbbm{E}_{x,r_{x}}[\Sigma_{2}]$
@@
This completes the proof.
IV-B Optimality of $\mathsf{NP}$-$\mathsf{SRPT}$ with finite task size
@@
$\mathbbm{E}{[F^{\pi^{*}}_{\rho}]}\geq\mathbbm{E}{[F^{\mathsf{SRPT}_{1,N}}_{\rho}]}.$ (11)
@@
$\lim_{n\rightarrow\infty}\frac{\mathbbm{E}{[F_{\rho}^{\mathsf{NP\text{-}SRPT}}]}-\mathbbm{E}[F^{\mathsf{SRPT}_{1,N}}_{\rho}]}{\mathbbm{E}[F^{\mathsf{SRPT}_{1,N}}_{\rho}]}=0,$ (12)
which holds according to Lemma 5 and inequality (IV-A)-(11).
V Optimality of $\mathsf{NP}$-$\mathsf{SRPT}$ Beyond Finite Task Size
Up to this point, we have focused on job size distributions with finite task, which is rather restrictive. It is natural to consider relaxations of this assumption. In this section, we turn to other classes of job size distributions with unbounded support on task workload. These results provide complement to our developments about the theory of the asymptotic optimality of $\mathsf{NP}$-$\mathsf{SRPT}$ .
V-A Warm up–random number of tasks
@@
Proof.
Consider an increasing convex function $g(x)$, by Jensen’s inequality, we have
@@
where $D(s)=\{s>0|\mathbbm{E}[e^{sp_{i}}]<\infty\}$. The proof is complete. ∎
V-B Optimality in $\mathsf{M}$$\mathsf{M}$$\mathsf{1}$
@@
$\mathsf{NP}$-$\mathsf{SRPT}$ is asymptotic optimal in $\mathsf{M}$$\mathsf{M}$$\mathsf{1}$.
@@
For $\mathsf{M}$$\mathsf{M}$$\mathsf{1}$ model and any work-conserving algorithm, let $n_{\mathrm{busy}}$ be the number of arrivals in a busy period, then
@@
Proof of Theorem 4:.
Firstly, it has been shown in [5] that, the average response time under SRPT in $\mathsf{M}$$\mathsf{M}$$\mathsf{1}$ satisfies
@@
$\mathbbm{E}{[F_{\rho}^{\mathsf{NP\text{-}SRPT}}]}-\mathbbm{E}{[F_{\rho}^{\mathsf{SRPT}_{1,N}}]}\leq O\Big(\log^{2}\frac{1}{1-\rho}\Big),$
which is a lower order term compared to $\mathbbm{E}[F^{\mathrm{SRPT-1}}_{\rho}]$. ∎
V-C Beyond Exponential Job Size Distribution
In the study of $\mathsf{NP}$-$\mathsf{SRPT}$, it is crucial to analyze the expected length of busy periods, which could help bound $\mathbbm{E}[\eta]$. Notably, previous works have established insights into the analysis of busy periods in $\mathsf{M}$$\mathsf{M}$$\mathsf{N}$ queuing systems[33, 2]. In this section, we extend the analysis in [33] to $\mathsf{M}$$\mathsf{G}$$\mathsf{N}$ and prove an upper bound on the expected length of the busy period under $\mathsf{NP}$-$\mathsf{SRPT}$, with the following assumption.
@@ V-C Beyond Exponential Job Size Distribution
Proof.
To analyze the busy periods under the $\mathsf{NP}$-$\mathsf{SRPT}$ policy, we classify two types of busy periods in an $N$-server system:
  • Full busy period. A time interval during which all $k\leq N$ servers are occupied. It starts when a new arrival finds k1k-1 customers already in the system and ends at the first departure epoch when exactly k1k-1 customers remain in the system. The length of such a busy period is denoted by $T_{k}$.
@@
$=$ $\mathbbm{P}(\text{departure first})\cdot\mathbbm{E}[\text{first departure time}|\text{departure first}]$
$+\mathbbm{P}(\text{arrival first})\cdot\mathbbm{E}[T_{k+1}|\text{start with}\{\mathsf{J}^{\prime}_{i}\}_{i=1}^{k}\cup\{\mathsf{J}_{k+1}\}]$
@@
Since we consider an $\mathsf{M}$$\mathsf{G}$ system, where job arrivals follow a Poisson process, the probability that an arrival occurs before a departure is:
$\mathbbm{P}(\text{arrival first})$
$=$ $\int_{0}^{\infty}\mathbbm{P}(\text{minimum remaining workload}$
$\text{of jobs under processing}\geq s)\cdot\lambda e^{-\lambda s}\mathrm{d}s$
@@
Using (V-C), we derive
$L_{k}(p_{\max})$
$\leq$ $\frac{1}{\mu}+\Big(1-\frac{1}{2}e^{-\frac{2r_{\max}^{2}}{N}}\Big)(L_{k+1}(p_{\max})+L_{k}(p_{\max}))$
$\Rightarrow$ $L_{k}(p_{\max})\leq\frac{2e^{\frac{2r_{\max}^{2}}{N}}}{\mu}+\Big(2e^{\frac{2r_{\max}^{2}}{N}}-1\Big)L_{k+1}(p_{\max}),$
and thus
$L_{k}(p_{\max})\leq\frac{2e^{\frac{2r_{\max}^{2}}{N}}}{\mu}+\Big(2e^{\frac{2r_{\max}^{2}}{N}}-1\Big)L_{k+1}(p_{\max}).$
@@
Let $\mathbb{E}[\mathsf{B}^{(N)}_{\mathsf{NP\text{-}SRPT}}$] denote the expected duration of a partial busy period, defined as the time from the first job arrival in an empty $N$-server system until the system becomes empty again under the $\mathsf{NP}$-$\mathsf{SRPT}$ policy. Then
$\mathbbm{E}[\mathsf{B}^{(N)}_{\mathsf{NP\text{-}SRPT}}$$\leq$ $\mathbbm{E}_{p_{\max}}\Big[\sum_{k=1}^{N}{L_{k}(p_{\max})}\Big]$
[…]
@@
[…]
This completes the proof, following the arguments in Proposition 1. ∎
@@
Let $f$ be a positive function defined in $0,\infty)$, the upper Matuszewska index is defined as the infimum of $\alpha$ for which there exists a constant $C=C(\alpha)$ such that for each $\bar{\lambda}>1$,
$\lim\nolimits_{x\rightarrow\infty}\frac{f(\lambda x)}{f(x)}\leq C\lambda^{\alpha},$
@@
In an $\mathsf{M}$$\mathsf{G}$/11 queue, if the upper Matuszewska index of the job size distribution is less than 2-2, then
@@
The average response time under $\mathsf{NP}$-$\mathsf{SRPT}$ is asymptotic optimal in $\mathsf{M}$$\mathsf{G}$$\mathsf{N}$ under Assumption 1, if upper Matuszewska index of job size distribution is less than 2-2 and
@@
V-D Discussions
In our analysis of $\mathsf{NP\text{-}SRPT}$, we consider a scenario where tasks within the same job cannot be executed in parallel, necessitating sequential processing. This setting naturally accommodates precedence constraints, such as those represented by a Directed Acyclic Graph (DAG), where tasks must follow a specific order (e.g., a topological sort). In this case, $\mathsf{NP\text{-}SRPT}$ remains effective by scheduling only the next ready task (with all predecessors completed) from the job with the shortest remaining processing time, respecting the topological order. This ensures that precedence constraints are satisfied while maintaining the greedy prioritization of $\mathsf{NP\text{-}SRPT}$. When tasks can be processed in parallel, finding the optimal allocation policy or analyzing the mean job response time for different allocation policies remains a challenging open problem [14].
VI Extensions to Scheduling with Unknown Job Sizes
The preceding analysis has established the asymptotic optimality of $\mathsf{NP}$-$\mathsf{SRPT}$ under the assumption that job sizes are known upon arrival. However, in many practical systems, this information is not available to the scheduler [38, 46]. This motivates the question of how to design effective scheduling policies when job sizes are unknown. In this section, we address this challenge by extending the analysis to unknown job size setting. We build upon the work of Scully et al. [38], and adapt their insights to the non-preemptive, multi-task environment.
We specifically consider non-preemptive version of the SOAP policy, monotone shortest expected remaining processing time ($\mathsf{M}$-$\mathsf{SERPT}$) and monotone Gittins ($\mathsf{M}$-$\mathsf{Gittins}$) policies, denoted as $\mathsf{NP}$-$\mathsf{M}$-$\mathsf{SERPT}_{N}$ and $\mathsf{NP}$-$\mathsf{M}$-$\mathsf{Gittins}_{N}$ respectively. The results for these policies apply to a wide range of job size distributions, categorized by their tail behavior. The key classes include O-Regularly Varying ($\mathsf{OR}$), Quasi-Decreasing Hazard Rate ($\mathsf{QDHR}$), and distributions in the Gumbel Domain of Attraction ($\mathsf{MDA}(\Lambda)$) [38]. For the sake of brevity and to maintain focus on the main results, the formal mathematical definitions for these classes are provided in Appendix Formal Definitions of Job Size Distribution Classes.
The following lemma relates the performance of these non-preemptive multi-server policies to their single-server preemptive counterparts in the heavy-traffic regime.
Lemma 7.
Let $\pi\in\{\mathsf{M\text{-}SERPT},\mathsf{M\text{-}Gittins}\}$ and consider a job size distribution in $\mathsf{OR}(-\infty,-1)\cup(\mathsf{QDHR}\cap\mathsf{MDA}(\Lambda))$, then
$\lim\nolimits_{\rho\to 1}\frac{\mathbbm{E}[F^{\mathsf{NP}\text{-}\pi_{N}}_{\rho}]}{\mathbbm{E}[F^{\pi_{1}}_{\rho}]}=1.$ (16)
Proof.
Since $\pi\in\{\mathsf{M\text{-}SERPT},\mathsf{M\text{-}Gittins}\}$ is a monotone non-decreasing SOAP policy, we can apply Lemma 1 to bound the difference in relevant work for job with size $x$ between the non-preemptive $N$-server policy $\mathsf{NP}\text{-}\pi$ and its preemptive single-server counterpart $\pi_{1,N}$:
$\mathsf{Rel}^{\mathsf{NP}\text{-}\pi}_{\leq x}(t)-\mathsf{Rel}^{\pi_{1,N}}_{\leq x}(t)\leq(N-1)(z^{\pi_{1,N}}_{x}+\eta),$ (17)
where $z_{x}^{\pi_{1,N}}$ represents the value of $z^{\pi_{1,N}}_{\mathsf{J}}(r^{\pi_{1,N}}(p_{\mathsf{J}},p_{\mathsf{J}}))$ for job $\mathsf{J}$ with size $x$. Following an argument analogous to that in [38], inequality (17) implies a bound on the expected response time:
$\mathbbm{E}[F^{\mathsf{NP}\text{-}\pi}_{\rho}]-\mathbbm{E}[F^{\pi_{1,N}}_{\rho}]$
$\leq$ $o(\mathbbm{E}[F^{\pi_{1,N}}_{\rho}])+O\left(\mathbbm{E}\left[\frac{\eta}{1-\rho(y^{\pi_{1,N}}_{x})}\right]\right),$ (18)
where $y^{\pi_{1,N}}_{x}$ denotes the new age cutoff. Let us define a threshold $\xi$ such that $\rho(y^{\pi_{1,N}}_{\xi})=\rho/2$, this implies:
$\frac{\rho}{2}=\int_{0}^{y^{\pi_{1,N}}_{\xi}}\lambda\bar{F}(t)\,\mathrm{d}t\leq\lambda y^{\pi_{1,N}}_{\xi}.$
As established in [38] for $\pi\in\{\mathsf{M\text{-}SERPT},\mathsf{M\text{-}Gittins}\}$, the age cutoff for the single-server policy, $y^{\pi_{1,N}}_{x}$, exhibits the following asymptotic behavior:
  • $y^{\pi_{1,N}}_{x}=\Theta(x)$ for job size distributions in $\mathsf{OR}(-\infty,-1)$.
  • $y^{\pi_{1,N}}_{x}=\Omega(x^{1/\gamma})$ for some $\gamma\geq 1$, for distributions in $\mathsf{QDHR}\cap\mathsf{MDA}(\Lambda)$.
In both scenarios, this yields a lower bound on $\xi$:
$\xi=\Omega\left(\left(\frac{\rho}{2\lambda}\right)^{\gamma}\right)$
for some constant $\gamma\geq 1$. Consequently, we can bound the final term in (VI) as,
$\mathbbm{E}\left[\frac{\eta}{1-\rho(y^{\pi_{1,N}}_{x})}\right]$
$\leq$ $\left(\frac{1}{1-\rho(y^{\pi_{1,N}}_{\xi})}+\frac{1}{\xi}\mathbbm{E}\left[\frac{x}{1-\rho(y^{\pi_{1,N}}_{x})}\right]\right)\cdot\mathbbm{E}[\eta]$
$\leq$ $O\left((1-\rho)^{-\varepsilon}\right)\cdot\mathbbm{E}[\eta],\quad\forall\varepsilon>0.$
Since $\mathbbm{E}[F^{\pi_{1,N}}_{\rho}]$ grows at least polynomially in $(1-\rho)^{-1}$, this error term is asymptotically negligible . Substituting this result into (VI) yields the desired limit. ∎
The lemma establishes that in the heavy-traffic limit, the mean response time under the non-preemptive, $N$-server policies $\mathsf{NP}$-$\mathsf{M}$-$\mathsf{SERPT}_{N}$ and $\mathsf{NP}$-$\mathsf{M}$-$\mathsf{Gittins}_{N}$ converges to that of their single-server, preemptive counterparts. This allows us to leverage known results for the single-server $\mathsf{M}$$\mathsf{G}$/11 setting. Specifically, the Gittins index policy is optimal, and it is known that $\mathsf{M}$-$\mathsf{Gittins}$ is asymptotically optimal (i.e., $\lim_{\rho\to 1}\mathbbm{E}[F^{\mathsf{M\text{-}Gittins}_{1}}_{\rho}]/\mathbbm{E}[F^{\mathsf{Gittins}_{1}}_{\rho}]=1$) and $\mathsf{M}$-$\mathsf{SERPT}$ is 2-competitive (i.e., $\mathbbm{E}[F^{\mathsf{M\text{-}SERPT}_{1}}_{\rho}]\leq 2\mathbbm{E}[F^{\mathsf{Gittins}_{1}}_{\rho}]$). These facts lead to our main theorem for the unknown job size case.
Theorem 6.
In the heavy-traffic limit ($\rho\to 1$) for a multi-task multi-server system with unknown job sizes from a distribution in $\mathsf{OR}(-\infty,-1)\cup(\mathsf{QDHR}\cap\mathsf{MDA}(\Lambda))$:
  1. (i)
    The $\mathsf{NP}$-$\mathsf{M}$-$\mathsf{Gittins}_{N}$ policy is asymptotically optimal.
  2. (ii)
    The $\mathsf{NP}$-$\mathsf{M}$-$\mathsf{SERPT}_{N}$ policy is asymptotically $2$-competitive.
VII Experimental Results